if the second digit of a 2 digit no 3 times of first digit and there position is changed and added to original no the answer is 88 wht are digits?
Answer by ankor@dixie-net.com(22740)

If the second digit of a 2 digit no 3 times of first digit and there position is changed and added to original no the answer is 88 what are digits?
:
Let x = the 1st digit
Let y = the 2nd digit
Then
10x + y = "the number"
:
Write an equation for each statement:
:
"if the second digit of a 2 digit no 3 times of first digit"
y = 3x
:
"their position is changed and added to original no the answer is 88"
10y+x + 10x+y = 88
11x + 11y = 88
:
Simplify divide by 8, results:
x + y = 8
:
what are digits?
:
Substitute 3x for y in the above equation
x + 3x = 8
4x = 8
x = 2 is the 1st digit
then
3(2) = 6 is the 2nd digit
:
26 is the number
;
;
Check solution 62 + 26 = 88